OpenStudy (anonymous):

Solve the equation 2^(x+2) = 16^x.

OpenStudy (anonymous):

2^(x+2)=2^4=16^x (x+2)=4 x=2

jhonyy9 (jhonyy9):

x+2=4x 2=3x x=2/3